# Interpreting GNN-based IDS Detections Using Provenance Graph Structural Features

### v5

@@ 4.1.2 Graph Structural Features
4.2 Surrogate Model Training
To obtain explanations from our security-aware graph structural features, we use them to train a global surrogate DT. By training a DT to agree with the predictions of a GNN model, we can interpret the DT to gain insights about the GNN’s decision-making process. To achieve the best agreement results, we enhance traditional DT training with data augmentation that iteratively increases the weight of incorrectly classified samples [trustee2022ccs].
We begin with a labelled set of graphs $D_{G}=(G,Y)$, which is used to train the GNN. GNN’s predictions on $D_{G}$ are collected, yielding $GNN(D_{G})=Y^{\prime}$. To prepare the dataset for training the DT, we extract the discriminative subgraph patterns and graph structural features (§ 4.1) $F$ and associate them with the GNN’s predictions $Y^{\prime}$ to create a labelled feature dataset $D_{F}=(F,Y^{\prime})$, which we split into train, validation, and testing to evaluate the surrogate DT.
Leveraging the methods of Jacobs et al. [jacobs2022ai], we use two-layer iterative dataset augmentation to train a series of DT models. At each iteration of the inner loop, all misclassified samples are duplicated to increase their weight in the next iteration; from this series of models, we select the one with the highest agreement among the DTs. This process is repeated several times in the outer loop, then the surrogate model with the highest mean agreement among those high-agreement DTs is selected as the final surrogate model for explanation.
4.3 Surrogate Model Interpretation
Once the surrogate DT is trained, we get instance-level explanation as it not only enables explanations based on local subgraph structures but also global structural changes. Each decision node within the DT represents a subgraph structure or a graph structural measure along with its threshold. Decision nodes closer to the root of the DT exert the greatest influence on the decision path; thus, we rank the nodes that participate in subgraphs based on their relative positioning in the decision tree. In cases where the decision path contains only graph structural measures without subgraph patterns, node-level explanations cannot be extractable, even if a DT tree is generated. While structural metrics alone face challenges in pinpointing specific nodes, they remain essential indicators of long-running persistent APT attacks. During an APT attack, though determined attackers may preserve node-level attributes, they typically cannot preserve benign program’s graph structure. Another scenario where explanations is invalid is when the surrogate model’s prediction disagrees with the GNN’s prediction.
System attributes (e.g., process/file names and socket IP/port) can be extracted in a post-processing step. Finally, the interpretable surrogate DT facilitates the construction of actionable, security-aware explanations for individual decisions. Since ProvExplainer produces a global surrogate DT, domain experts can use it to deepen their understanding of the GNN’s decision-making process. In our experiments, we found that limiting DT nodes to a depth of 4 provided the most effective explanations in our datasets. Lower depths failed to capture shapes needed for complex APT scenarios, while higher depths introduced irrelevant system behaviors. Due to the DT’s susceptibility to overfitting, restricting the depth was necessary.
5 Evaluation
@@
    Explanation Accuracy. Can ProvExplainer explain APT and Fileless malware detection (§ 5.3,§ 5.4)?
@@
    Comparison with SOTA GNN Explainers. How do the explanations of ProvExplainer compare against those of SOTA GNN explainers (GNNExplainer [ying2019gnnexplainer], PGExplainer [luo2020parameterized], and SubgraphX [yuan2021explainability]) (§ 5.5)?
5.1 Evaluation Protocols
@@
We evaluate ProvExplainer against different APT detection datasets and Fileless Malware  datasets: (1) publicly available APT attack simulations [darpa], (2) APT dataset from a recent study [mukherjee2023sec], and (3) execution traces of Fileless Malware [survivialism2021sp]. We implemented two general purpose SOTA GNN-based detector models: GAT  [velivckovic2017graph] and GraphSAGE  [hamilton2017inductive], following the approach of recent explanation literature [herath2022cfgexplainer, kosan2023gnnx]. Recent GNN based anomaly detection systems [zengy2022shadewatcher, rehman2024flash, jia2024magic, cheng2024kairos, goyal2024rcaid] rely on custom node and edge embeddings for security tasks and they do not use the GNN for anomaly detection task but rather use them for extracting graph embeddings and use another downstream model for anomaly detection, so we did not evaluate against these specialized solutions.
@@
We conducted an ablation study to understand the impact of different feature sets consisting of subgraph patterns and graph structural features. Finally, we evaluated the explanations given by ProvExplainer against those of SOTA GNN explainers [ying2019gnnexplainer, luo2020parameterized, yuan2021explainability] In § 6, we present case studies highlighting specific instances of APT detection and offer in-depth insights into how our approach provides security relevant explanations.
@@
Evaluation Metrics. In our evaluation of ProvExplainer, we focus on three critical aspects. The first is the agreement of the surrogate DTs with the GNN model, which we measure using the weighted macro averaged (WMA) F1 score of the surrogate DT’s predictions with respect to the GNN’s predictions. The agreement metric gauges the faithfulness of the DT in replicating the conclusions of GNNs. The choice of the WMA F1 score accounts for the data imbalance issue in the anomaly detection datasets.
The second metric is $\text{fidelity}^{+}$  and $\text{fidelity}^{-}$  used by different GNN explainers [yuan2020explainability] which measures if we remove the important subgraph pattern does the class label flip and if we remove unimportant subgraph patterns the class label should not change. A good explainer is the one that will have a high $\text{fidelity}^{+}$  and a low $\text{fidelity}^{-}$
@@
To evaluate ProvExplainer’s and SOTA GNN explainers’ proficiency in identifying security-relevant entities, we define precision and recall metrics with respect to documented entities § 2.4. A graph explanation method $\mathcal{E}$ will yield a total ordering over $\mathcal{V}$ for a given graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ and graph model $\mathcal{M}$. Let $\mathcal{D}$ be the set of documented entities. Precision is the proportion of explanation nodes that are documented, and recall is the fraction of documented entities retrieved: $precision(V_{k},k,\mathcal{D})=\frac{|V_{k}\cap\mathcal{D}|}{k}$, and $recall(V_{k},\mathcal{D})=\frac{|V_{k}\cap\mathcal{D}|}{|\mathcal{D}|}$.
5.2 Evaluation Tasks
@@
APT Detection. We utilized the DARPA Transparent Computing (TC) Data Releases [darpa] and previous literature [mukherjee2023sec] for APT attack detection. This dataset, encompassing various OSes, provides a comprehensive basis for advanced security research. The DARPA APTs were designed to attack a system which consists of long-running processes and captured the stealthy attack vectors frequently employed by advanced adversaries. We particularly focused on three DARPA datasets used by previous studies [zengy2022shadewatcher, cheng2024kairos, rehman2024flash, goyal2024rcaid]: FiveDirections, Trace, and Theia. However, these tasks were conducted in a simulated environment and only lasted for two weeks, involving a limited number of hosts. Therefore, we also evaluated against the APT scenarios (e.g., Enterprise APT and Supply-Chain APT) conducted by  [mukherjee2023sec], which were performed with realistic benign background workloads. Interested readers can find detailed statistics presented in Table IV in the appendix.
@@
Fileless Malware Detection. For Fileless Malware detection, we targeted a family of stealthy malware samples that impersonate benign programs, evading conventional security solutions but are detectable by GNN-based provenance analysis. The malware samples were chosen in accordance with guidelines from the literature [kuchler2021does, avllazagaj2021malware] to minimize experimental bias and ensure freshness. The Fileless Malware dataset includes various categories [kuchler2021does], including banking Trojans, spyware, ransomware, and malware installers; detailed statistics presented in Table IV in the appendix.
@@
The Banking banking trojan [kuchler2021does] steals banking credentials from victim machines and spreads through spam and compromised download links. To camouflage interactions with .dlls and temporary files, it masquerades as svchost.exe. The Spyware, Ulise [kuchler2021does], is a multi-purpose Trojan that can establish remote access connections, capture keyboard input, collect system information, download/upload files, drop other malware into the infected system, and perform encryption. Because it propagates through the network sockets and interacts with many system files, it masquerades as python.exe as it needs a target that exhibits diverse behavioral patterns.
Table II: Surrogate DTs have high agreement ($\geq$ 0.85) with the decisions of GNN measured using the WMA F1 score.
Dataset GAT Surrogate DT (agree w/ GAT) GraphSAGE Surrogate DT (agree w/ GraphSAGE)
APT Dataset from [mukherjee2023sec]
Enterprise 0.82 0.95 0.81 0.94
Supply-Chain 0.82 0.94 0.80 0.97
Average 0.82 0.95 0.81 0.95
@@
Fileless Malware from [survivialism2021sp]
@@
5.3 Graph Structural Feature Evaluation
To answer RQ1, Table II demonstrates the effectiveness of surrogate DTs in mirroring the decision process of GNN models like GAT  and GraphSAGE . Agreement between the surrogate DT and the GNN is an important metric for two reasons: (1) surrogate explanations are only valid when the surrogate agrees with the GNN, and (2) high agreement indicates that the surrogate model is a good approximation of the GNN’s decision-making process.
ProvExplainer exhibits the highest agreement($>$ 97%) with the GNN models on the Fileless Malware  datasets, but the APT datasets only showed good agreement ($>$ 93%). The surrogate DTs’ superior performance in Fileless Malware  datasets can be attributed to ProvExplainer’s ability to capture local subgraph patterns with high efficiency but APT attacks that are long running and utilize evasive tactics (noted by previous studies [mukherjee2023sec, survivialism2021sp]) are hard to detect and explain. The dataset such as FiveDirections and conhost.exe are dominated by attackers using stealthy techniques such as living-off-the-land, which involve memory object interactions which are currently not captured in our provenance graphs. The absence of the distinguishing features impairs the effectiveness of ProvExplainer.
We observed that GAT  outperforms GraphSAGE  in majority of the datasets. This trend is attributed to GAT ’s ability to assign varying importance to different neighborhood structures and capture long-range structural dependencies through its attention mechanism. Conversely, GraphSAGE , which samples a fixed-size neighborhood, overlooks subtle structural changes, such as those resulting from elusive APT attacks. ProvExplainer’s graph structural features enable surrogate DTs to efficiently approximate GNN models’ decision-making process on in-distribution data. Although the current features are sensitive to attacker’s TTP, ProvExplainer is expandable to incorporate additional features to extend support to for new TTP explanation.
@@ 5.3 Graph Structural Feature Evaluation
Table III: Agreement (higher is better) measured using WMA F1 score of surrogate DTs with the GAT  model across different features.
Dataset Top-$K$ Subgraphs All Top 5 Subgraphs Graph Structural Features All Graph Structural Feat. All Features
1 2 3 4 5 Deg. Cen. Clos. Cen. Bet. Cen Eig. Cen. Clus. Coeff. Clus. Tri.
APT Dataset from [mukherjee2023sec]
Enterprise 0.90 (-0.05) 0.90 (-0.05) 0.90 (-0.05) 0.90 (-0.05) 0.90 (-0.05) 0.92 (-0.03) 0.92 (-0.03) 0.92 (-0.03) 0.92 (-0.03) 0.95 (-0.00) 0.95 (-0.00) 0.95 (-0.00) 0.82 (-0.13) 0.95
Supply-Chain 0.79 (-0.16) 0.85 (-0.09) 0.86 (-0.08) 0.87 (-0.07) 0.87 (-0.07) 0.93 (-0.01) 0.93 (-0.01) 0.93 (-0.01) 0.94 (-0.00) 0.94 (-0.00) 0.94 (-0.00) 0.94 (-0.00) 0.82 (-0.12) 0.94
Average 0.84 (-0.10) 0.88 (-0.07) 0.88 (-0.07) 0.89 (-0.06) 0.89 (-0.06) 0.93 (-0.02) 0.93 (-0.02) 0.93 (-0.02) 0.93 (-0.02) 0.95 (-0.00) 0.95 (-0.00) 0.95 (-0.00) 0.82 (-0.12) 0.95
DARPA APT Dataset [darpa]
FiveDirections 0.72 (-0.16) 0.75 (-0.13) 0.75 (-0.13) 0.72 (-0.16) 0.72 (-0.16) 0.72 (-0.16) 0.72 (-0.16) 0.72 (-0.16) 0.78 (-0.10) 0.78 (-0.10) 0.78 (-0.10) 0.84 (-0.04) 0.86 (-0.02) 0.88
Trace 0.93 (-0.00) 0.93 (-0.00) 0.93 (-0.00) 0.93 (-0.00) 0.87 (-0.07) 0.87 (-0.07) 0.87 (-0.07) 0.87 (-0.07) 0.87 (-0.07) 0.93 (-0.00) 0.93 (-0.00) 0.93 (-0.00) 0.93 (-0.00) 0.93
Theia 0.83 (-0.17) 0.83 (-0.17) 0.89 (-0.11) 0.89 (-0.11) 0.89 (-0.11) 0.89 (-0.11) 1.00 (-0.00) 1.00 (-0.00) 1.00 (-0.00) 1.00 (-0.00) 1.00 (-0.00) 1.00 (-0.00) 1.00 (-0.00) 1.00
Average 0.86(-0.11) 0.87(-0.09) 0.88(-0.08) 0.88(-0.08) 0.88(-0.08) 0.91(-0.06) 0.92(-0.05) 0.92(-0.05) 0.95(-0.01) 0.96(-0.01) 0.96(-0.01) 0.96(-0.01) 0.94(-0.03) 0.97
Fileless Malware from [survivialism2021sp]
wscript.exe 0.93 (-0.07) 0.93 (-0.07) 0.93 (-0.07) 1.00 (-0.00) 1.00 (-0.00) 0.98 (-0.02) 0.98 (-0.02) 0.98 (-0.02) 1.00 (-0.00) 1.00 (-0.00) 1.00 (-0.00) 1.00 (-0.00) 1.00 (-0.00) 1.00
cscript.exe 0.86 (-0.14) 0.86 (-0.14) 0.93 (-0.07) 0.93 (-0.07) 1.00 (-0.00) 1.00 (-0.00) 1.00 (-0.00) 1.00 (-0.00) 1.00 (-0.00) 1.00 (-0.00) 1.00 (-0.00) 1.00 (-0.00) 1.00 (-0.00) 1.00
reg.exe 0.92 (-0.08) 0.92 (-0.08) 0.92 (-0.08) 0.92 (-0.08) 0.92 (-0.08) 0.92 (-0.08) 0.92 (-0.08) 0.92 (-0.08) 1.00 (-0.00) 1.00 (-0.00) 1.00 (-0.00) 1.00 (-0.00) 1.00 (-0.00) 1.00
python.exe 0.59 (-0.41) 0.66 (-0.34) 0.66 (-0.34) 0.66 (-0.33) 0.65 (-0.35) 0.94 (-0.06) 0.94 (-0.06) 0.94 (-0.06) 0.99 (-0.00) 0.99 (-0.00) 0.99 (-0.00) 0.99 (-0.00) 0.99 (-0.00) 0.99
explore.exer 0.74 (-0.25) 0.91 (-0.07) 0.91 (-0.07) 0.91 (-0.07) 0.91 (-0.07) 0.91 (-0.07) 0.91 (-0.07) 0.91 (-0.07) 0.98 (-0.00) 0.98 (-0.00) 0.98 (-0.00) 0.98 (-0.00) 0.98 (-0.00) 0.98
netsh.exe 0.93 (-0.04) 0.91 (-0.07) 0.91 (-0.07) 0.89 (-0.09) 0.89 (-0.09) 0.89 (-0.09) 0.89 (-0.09) 0.91 (-0.07) 0.98 (-0.00) 0.98 (-0.00) 0.98 (-0.00) 0.98 (-0.00) 0.98 (-0.00) 0.98
net.exe 0.94 (-0.02) 0.94 (-0.02) 0.94 (-0.02) 0.94 (-0.02) 0.94 (-0.02) 0.91 (-0.06) 0.91 (-0.06) 0.91 (-0.06) 0.97 (-0.00) 0.97 (-0.00) 0.97 (-0.00) 0.97 (-0.00) 0.97 (-0.00) 0.97
rundll32.exe 0.85 (-0.10) 0.85 (-0.10) 0.85 (-0.10) 0.85 (-0.10) 0.85 (-0.10) 0.85 (-0.10) 0.85 (-0.10) 0.85 (-0.10) 0.96 (-0.00) 0.96 (-0.00) 0.96 (-0.00) 0.96 (-0.00) 0.96 (-0.00) 0.96
mshta.exe 0.75 (-0.20) 0.75 (-0.20) 0.75 (-0.20) 0.75 (-0.20) 0.75 (-0.20) 0.90 (-0.05) 0.90 (-0.05) 0.90 (-0.05) 0.95 (-0.00) 0.95 (-0.00) 0.95 (-0.00) 0.95 (-0.00) 0.95 (-0.00) 0.95
schtasks.exe 0.98 (-0.02) 0.98 (-0.02) 0.98 (-0.02) 0.98 (-0.02) 0.98 (-0.02) 0.98 (-0.02) 1.00 (-0.00) 1.00 (-0.00) 1.00 (-0.00) 1.00 (-0.00) 1.00 (-0.00) 1.00 (-0.00) 0.93 (-0.07) 1.00
conhost.exe 0.91 (-0.02) 0.89 (-0.00) 0.89 (-0.00) 0.89 (-0.00) 0.89 (-0.00) 0.89 (-0.00) 0.93 (-0.04) 0.89 (-0.00) 0.89 (-0.00) 0.89 (-0.00) 0.89 (-0.00) 0.89 (-0.00) 0.89 (-0.00) 0.89
svchost.exe 0.98 (-0.02) 0.98 (-0.02) 0.98 (-0.02) 0.98 (-0.02) 0.98 (-0.02) 0.98 (-0.02) 0.98 (-0.02) 0.98 (-0.02) 1.00 (-0.00) 1.00 (-0.00) 1.00 (-0.00) 1.00 (-0.00) 0.86 (-0.14) 1.00
Average 0.86 (-0.11) 0.88 (-0.09) 0.89 (-0.09) 0.89 (-0.08) 0.90 (-0.08) 0.93 (-0.05) 0.93 (-0.04) 0.93 (-0.04) 0.98 (-0.00) 0.98 (-0.00) 0.98 (-0.00) 0.98 (-0.00) 0.96 (-0.02) 0.98
5.4 Ablation Study
Table III shows the contributions of different features to the overall agreement of surrogate DTs approximating a GAT  model. Interested readers may find this study with GraphSAGE  in the appendix, Table V. We separate the features in two groups, subgraphs and graph structural features. We evaluate each group individually and then we look at the their combined effect. We further separate the group into individual features and look at the effect of each constituting features in the group. In the subgraph group, we use the top-$K$ subgraphs as features, considering each of them one by one, then followed by using all top-$K$ subgraphs. Second, we evaluate different graph structural features individually and then collectively.
The top-$K$ subgraphs demonstrate varying levels of agreement across the datasets, in the APT dataset, we consider each of the top-5 subgraphs individually shows different level of agreement for both Enterprise and Supply-Chain dataset, indicating that different subgraphs captures different signals in the graph to explain, therefore the alignment with the original model predictions stays constant. The trend is similar for the DARPA TC  and Fileless Malware  datasets, although the increase in agreement levels tends to plateau as $K$ approaches 4, suggesting that the initial few subgraphs capture the majority of informative patterns. Dissimilar to subgraph patterns when we examining individual structural features the performance varies significantly depending on the feature used. For instance, eigenvector centrality and clustering coefficient consistently shows higher agreement in the DARPA TC  dataset, particularly in Trace and Theia both achieving agreement scores of 1.00. In contrast, closeness and betweenness centrality features are comparatively lower in agreement for most datasets, possibly indicating that these features does not capture unique graph characteristics that are distinguishable.
Combining the top-5 subgraphs and all structural features generally results in high alignment with the model, achieving around 0.94 in agreement for Fileless Malware datasets. An interesting pattern is that the Fileless Malware  dataset consistently performs exceptionally well, particularly with individual features like eigenvector centrality and clustering triangles, which achieve near-perfect alignment scores. This suggests that the structural properties in the Fileless Malware dataset are distinctively different compared to their benign programs they are masquerading. In APT dataset where the attacker uses stealthy techniques more holistic feature combinations are needed for optimal performance. This trend suggests that when subgraph structures are combined with centrality and clustering metrics, they can capture a comprehensive representation of the graph structure, effectively enhancing the model’s agreement. The high performance across all datasets with this combination indicates that a more holistic representation of both the subgraph and graph structural features offers better alignment than individual metrics.
5.5 ProvExplainer vs. SOTA Explainers

### v7

@@ V-A Feature Extraction
V-B Surrogate Model Training
To obtain explanations from our security-aware graph structural features, we use them to train a global surrogate DT. By training a DT to agree with the predictions of a GNN model, we can interpret the DT to gain insights about the GNN’s decision-making process. To achieve the best agreement results, we enhance traditional DT training with data augmentation that iteratively increases the weight of incorrectly classified samples [61].
We begin with a labeled set of graphs $D_{G}=(G,Y)$, which is used to train the GNN. GNN’s predictions on $D_{G}$ are collected, yielding $GNN(D_{G})=Y^{\prime}$. To prepare the dataset for training the DT, we extract the discriminative subgraph patterns and graph structural features (§ V-A) $F$ and associate them with the GNN’s predictions $Y^{\prime}$ to create a labelled feature dataset $D_{F}=(F,Y^{\prime})$, which we split into train, validation, and testing to evaluate the surrogate DT.
V-C Surrogate Model Interpretation
We obtain instance-level explanations after training the surrogate DT. Each decision node in the DT corresponds either to a discriminant subgraph pattern or to a graph-structural measure (with its associated threshold). Nodes closer to the root of the tree exert greater influence on the decision path; therefore, we rank subgraph-participating nodes based on their relative position in the DT.
In cases where (1) the surrogate model’s prediction diverges from that of the GNN, and (2) the decision path is composed solely of structural measures without subgraph patterns, node-level explanations cannot be extracted, even though a DT is generated. While structural metrics alone cannot precisely localize specific nodes, they remain crucial indicators of long-running, persistent APT campaigns. This is because during an APT attack, adversaries may preserve node-level attributes, but they cannot maintain the benign program’s graph structure.
Finally, the interpretable surrogate DT enables the construction of actionable, security-aware explanations for individual predictions. Using ProvExplainer’s global surrogate DT and participating node’s system attributes (e.g., process names, file names, and IP/PORT), domain experts can leverage this enriched contextual information to understand the decision-making process of the GNN.
VI Evaluation
@@
    Explanation Accuracy. Can ProvExplainer accurately explain APT and Fileless Malware detections (§ VI-D and § VI-E)?
@@
    Comparison with SOTA GNN Explainers. How do ProvExplainer’s explanations compare against SOTA GNN explainers (e.g., GNNExplainer [31], PGExplainer [32], and SubgraphX [33]) (§ VI-F)?
  • RQ3:
    Explainability Effectiveness. How much actionable are ProvExplainer’s explanations (§ VI-G)?
VI-A Evaluation Protocols
@@
We evaluate ProvExplainer on a different APT detection and Fileless Malware datasets: (1) publicly available APT attack simulations [darpa], (2) an APT dataset from a recent study [46], and (3) execution traces of Fileless Malware [47]. We implement two general-purpose SOTA GNN-based detector models, GAT [68] and GraphSAGE [69], following the methodology adopted in recent explanation literature [60, 62]. Our focus is to first thoroughly understand and explain IDS that directly leverage GNNs for security-critical tasks, before addressing more complex IDS architectures that use GNNs to extract graph embeddings for downstream anomaly detection, as discussed in [26]. Accordingly, we do not compare against recent GNN-based IDS [56, 20, 25, 55, 19, 21, 27, 26] that rely on custom graph embeddings and employ a separate model for final detection.
@@
We conducted an ablation study to understand the impact of different feature sets (e.g., subgraph patterns and graph structural features). We also evaluate the explanations given by ProvExplainer against SOTA GNN explainers [31, 32, 33]. Finally, we conduct a non-trivial human-centered evaluation to validate the interpretability of the explanations from a security-analyst perspective. In § VII, we present case studies explaining APT detection and offer in-depth insights into how ProvExplainer provide security-relevant explanations.
VI-B Evaluation Metrics.
@@
In our evaluation of ProvExplainer, we focus on three critical aspects. The first is the agreement of the surrogate DTs with the GNN model, which we measure using the weighted macro averaged (WMA) F1 score of the surrogate DT’s predictions with respect to the GNN’s predictions. The agreement metric gauges the faithfulness of the DT in replicating the conclusions of GNNs. WMA F1 score helps to account for the data imbalance issue present in the anomaly detection datasets.
[…]
@@
To evaluate ProvExplainer’s and SOTA GNN explainers’ proficiency in identifying security-relevant entities, we define precision and recall metrics with respect to documented entities § II-D. Let $\mathcal{D}$ be the set of documented entities and using the above definitions for $\mathcal{V}_{i}$ and $k$. Precision is the proportion of explanation nodes that are documented, and recall is the fraction of documented entities retrieved: $precision(V_{k},k,\mathcal{D})=\frac{|V_{k}\cap\mathcal{D}|}{k}$, and $recall(V_{k},\mathcal{D})=\frac{|V_{k}\cap\mathcal{D}|}{|\mathcal{D}|}$.
Lastly, to assess the actionability of the explanations from a security analyst’s perspective, we introduce a graph traversal distance metric, where we compute the shortest-path distance between each node recommended by an explainer and its corresponding ground-truth node (§ II-D). Let $\mathcal{V}_{\text{GT}}$ denote the set of ground-truth nodes and $\mathcal{V}_{\text{expl}}$ the set of source nodes highlighted by the explainer. The distance metric is defined as: $distance_{traversal}=\frac{1}{|\mathcal{V}_{\text{GT}}|}\sum_{i=1}^{|\mathcal{V}_{\text{GT}}|}d(v^{\text{expl}}_{i},v^{\text{GT}}_{i})$, where $d(v^{\text{expl}}_{i},v^{\text{GT}}_{i})$ is the shortest-path node distance between the $i$-th recommended node and its matched ground-truth node in the graph. This node-based distance is then normalized by the number of ground-truth nodes $|\mathcal{V}_{\text{GT}}|$, ensuring a consistent scale across graphs. The intuition behind this metric is that a lower distance indicates the explainer’s recommendation is closer to the true malicious entities, thereby reducing the analyst’s cognitive and temporal burden during an investigation.
(a) DARPA APT Dataset [darpa]
(b) APT Dataset from [46]
(c) Fileless Malware from [47]
Figure 3: Agreement (higher is better) measured using WMA F1 of surrogate decision trees versus the GAT model across different feature sets.
VI-C Evaluation Tasks
@@
APT Detection. We utilized the DARPA Transparent Computing (TC) Data Releases [darpa] and previous literature [46] for APT attack detection. This dataset, encompassing various OSes, provides a comprehensive basis for advanced security research. The DARPA APTs were designed to attack a system which consists of long-running processes and captured the stealthy attack vectors frequently employed by advanced adversaries. We particularly focused on three DARPA datasets used by previous studies [56, 19, 20, 21]: FiveDirections, Trace, and Theia. However, these tasks were conducted in a simulated environment and only lasted for two weeks, involving a limited number of hosts. Therefore, we also evaluated against the APT scenarios (e.g., Enterprise APT and Supply-Chain APT) conducted by [46], which were performed with realistic benign background workloads. Interested readers can find detailed statistics presented in Table IV.
@@
Fileless Malware Detection. For Fileless Malware detection, we targeted a family of stealthy malware samples that impersonate benign programs, evading conventional security solutions but are detectable by GNN-based provenance analysis. The malware samples were chosen in accordance with guidelines from the literature [70, 71] to minimize experimental bias and ensure freshness. The Fileless Malware dataset includes various categories [70], including banking Trojans, spyware, ransomware, and malware installers; detailed statistics presented in Table IV in the appendix.
@@
The Banking banking trojan [70] steals banking credentials from victim machines and spreads through spam and compromised download links. To camouflage interactions with .dlls and temporary files, it masquerades as svchost.exe. The Spyware, Ulise [70], is a multi-purpose Trojan that can establish remote access connections, capture keyboard input, collect system information, upload files, drop other malware into the infected system, and perform encryption. Because it propagates through the network sockets and interacts with many system files, it masquerades as python.exe as it needs a target that exhibits diverse behavioral patterns.
Dataset
GAT
Surrogate DT
(agree w/ GAT)
GraphSAGE
Surrogate DT
(agree w/ GraphSAGE)
@@
APT Dataset from [46]
Enterprise 0.82 0.95 0.81 0.94
Supply-Chain 0.82 0.94 0.80 0.97
Average 0.82 0.95 0.81 0.95
Fileless Malware from [47]
@@
Table I: Surrogate DTs have high agreement ($\geq$ 0.85) with the decisions of GNN measured using the WMA F1 score.
VI-D Graph Structural Feature Evaluation
To answer RQ1, Table I demonstrates the effectiveness of surrogate DTs in mirroring the decision process of GNN models like GAT and GraphSAGE . Agreement between the surrogate DT and the GNN is an important metric for two reasons: (1) surrogate explanations are only valid when the surrogate agrees with the GNN, and (2) high agreement indicates that the surrogate model is a good approximation of the GNN’s decision-making process.
ProvExplainer exhibits the highest agreement(>> 97%) with the GNN models on the Fileless Malware datasets, but the APT datasets only showed good agreement (>> 93%). The surrogate DTs’ superior performance in Fileless Malware datasets can be attributed to ProvExplainer’s ability to capture local subgraph patterns with high efficiency but APT attacks that are long running and utilize evasive tactics (noted by previous studies [46, 47]) are harder to detect and explain. The dataset such as FiveDirections and conhost.exe are dominated by attackers using stealthy techniques such as living-off-the-land, which involve memory object interactions that are currently not captured in our provenance graphs.
The absence of the distinguishing features or rare features that are exclusively present in the testing dataset impairs the effectiveness of ProvExplainer. We observed that GAT outperforms GraphSAGE in majority of the datasets. This trend is attributed to GAT ’s ability to assign varying importance to different neighborhood structures and capture long-range structural dependencies through its attention mechanism. Conversely, GraphSAGE , which samples a fixed-size neighborhood, overlooks subtle structural changes, such as those resulting from elusive APT attacks. ProvExplainer’s graph structural features enable surrogate DTs to efficiently approximate GNN models’ decision-making process on in-distribution data. Although the current features are sensitive to attacker’s TTPs, ProvExplainer is expandable to incorporate new features to extend support for the new TTPs.